OpenStudy (anonymous):

What are the intercepts of the equation? Graph the equation. 5x + 2y = 10

OpenStudy (anonymous):

a)x-intercept: (5,0) y-intecept: (0,2) b)x-intercept: (2,0) y-intercept: (0,5) c)x-intercept: (2,0) y-intecept: (0,5) d)x-intercept: (5,0) y-intercept: (0,2)

hartnn (hartnn):

To get x-intercept put y=0 in the equation 5x + 2y = 10 to get y-intercept put x=0 in the equation 5x + 2y = 10

hartnn (hartnn):

what u get when u put y=0 in the equation 5x + 2y = 10?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

hold on :p

OpenStudy (anonymous):

so 2 @hartnn

hartnn (hartnn):

yes, x-intercept is 2 (2,0) what about y intercept ?

OpenStudy (anonymous):

so (2,5)?:)

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

you get y intercepts as 5 so, (0,5) so, 2 intercepts are (2,0) and (0,5)

OpenStudy (anonymous):

Okay, thanks so much for explaining <3 @hartnn